OpenStudy (akohl):

What is the intensity level of a sound with an intensity of 0.000127 W/m2? 120 dB 65 dB 81 dB 72 dB @Abhisar

OpenStudy (akohl):

@aaronq

OpenStudy (aaronq):

\(\sf \beta=10dB*10^{\dfrac{I}{I_o}}\) \(\sf I_o=1.0*10^{-12}~W/m^2\)

OpenStudy (akohl):

Im getting an insanely high number

OpenStudy (aaronq):

oh sorry, i screwed up the formula its: \(\sf \beta =10dB*log(\dfrac{I}{I_o})\)